Observers:  Luke Harvey, Simon Prentice

Support: Maxime Deckers, Mark Magee

TAT: Steve Schulze, Jacob Teffs, Santiago Gonzalez

EFOSC Calibrations: bias, calib, UBVRI flats, gri flats


Start: humidity 30, wind 8.7, seeing 1.4

Middle: humidity 20, wind 12.6, seeing 1.2

End: humidity 6, wind 12.8, seeing -


Clouds from 01:15 - 02:50.

Clouds from 03:30 - 05:40.

TO P2 not updating, unable to add OBs.

Wind caused pointing restrictions from 5:40.

The seeing monitor was not operational for the latter half of the night.

Started with EFOSC